# Editing Across Languages: A Survey of Multilingual Knowledge Editing

### v1

@@ 5.2 Cross Lingual Transferability
The strongest predictor of successful cross-lingual propagation is the degree of linguistic similarity between the source and target languages. Edits propagate better within the same language family (e.g., English to Spanish or German (see Figure 2)) than between distant language families (e.g., English to Chinese or Arabic) Zhang et al. (2025); Wei et al. (2025); Nie et al. (2025). Some studies further suggest that shared writing systems or similar orthographic structures (e.g., Latin-based scripts) may offer marginal additional benefits to transferability Nie et al. (2025); Xu et al. (2023).  Beniwal et al. (2024) also showed that script differences (e.g., Devanagari vs. Tamil) reduce edit transferability, highlighting script divergence as a secondary but relevant limiting factor. Recent work further underscores the role of subject token similarity in transfer success:  Tamayo et al. (2024) showed that edits are more likely to propagate when the embedding representations of subject entities are closely aligned across languages.
@@
The language in which the edit is applied has a notable impact on cross-lingual propagation. Edits performed in high-resource languages such as English generally transfer more effectively to other languages than edits made in lower-resource or non-Latin languages such as Chinese Wu et al. (2025); Zhang et al. (2025); Xie et al. (2024); Nie et al. (2025). This asymmetry likely reflects the dominant role of English and other high-resource languages during pretraining and fine-tuning of multilingual language models.
@@ 5.3 Impact of Model Size and Capability
Figure 3: Generality scores for BLOOMZ and LLaMA models across increasing model sizes. ReMaKE Wang et al. (2024c) (BLOOMZ) results are extracted for mzsRE generality, averaged across 10 languages. IKE Nie et al. (2025) (LLaMA) results are extracted for mzsRE generality under zero-shot settings, averaged across 53 languages. Results show that generality consistently improves with model scale for both models.
@@ 6 Discussion
Multilingual knowledge editing (MKE) has emerged as a promising extension of the model editing paradigm, enabling factual updates that generalize across languages. While recent progress has introduced diverse method families, no single method fully solves the core challenge of reliable multilingual propagation. Parameter-based methods are limited by representation anisotropy; memory-based methods face retrieval bottlenecks and storage overhead; prompting is unreliable and fragile across languages; fine-tuning is static and prone to interference; and hypernetwork approaches, though promising, remain underexplored. These limitations motivate several critical open challenges.
@@ 6.1 Open Challenges
Multilingual LLMs often encode languages into distinct internal subspaces, creating structural anisotropy that inhibits cross-lingual transfer of factual edits. Bridging this gap remains the most fundamental unsolved obstacle in MKE.
@@
Most MKE studies evaluate only high-resource languages (e.g., English, Chinese). Systematic evaluation on low-resource, morphologically complex, and under-represented languages (e.g., Swahili, Hebrew) remains largely absent.
@@
The field still lacks unified evaluation protocols and benchmarks. While MLaKE, Bi-ZsRE, and MZSRE are important steps forward, inconsistencies in construction, metrics, and language coverage persist. Many rely on machine translation outputs, introducing noise. Low-resource and morphologically rich languages remain underrepresented.
@@ 6.2 Opportunities
Development of Robust Multilingual Benchmarks
@@ Appendix A Details on MKE Benchmarks
MzSRE Wang et al. (2024c) is a machine-translated version of the ZsRE dataset in 12 languages: English, Czech, German, Dutch, Spanish, French, Portuguese, Russian, Thai, Turkish, Vietnamese, and Chinese. Translations were generated using Google Translate without human revision. Each sample is used to evaluate reliability, generality, locality, and portability.
@@
Eva-KELLM Wu et al. (2023) extends knowledge editing evaluation to full documents. Starting from the Counterfact dataset Meng et al. (2023a), ChatGPT was used to generate documents corresponding to each edited fact. The dataset consists of 8,882 English documents and 6,930 Chinese documents.

### v2

@@ 5.2 Cross-Lingual Transferability
The strongest predictor of successful cross-lingual propagation is the degree of linguistic similarity between the source and target languages. Edits propagate better within the same language family (e.g., English to Spanish or German (see Figure 2)) than between distant language families (e.g., English to Chinese or Arabic) Mousi et al. (2025); Zhang et al. (2025); Wei et al. (2025). Some studies further suggest that shared writing systems or similar orthographic structures (e.g., Latin-based scripts) may offer marginal additional benefits to transferability Nie et al. (2025); Xu et al. (2023).  Beniwal et al. (2024) also showed that script differences (e.g., Devanagari vs. Tamil) reduce edit transferability, highlighting script divergence as a secondary but relevant limiting factor. Recent work further underscores the role of subject token similarity in transfer success:  Tamayo et al. (2024) showed that edits are more likely to propagate when the embedding representations of subject entities are closely aligned across languages.
@@
The language in which the edit is applied has a notable impact on cross-lingual propagation. Edits performed in high-resource languages such as English generally transfer more effectively to other languages than edits made in lower-resource or non-Latin languages such as Arabic Mousi et al. (2025) or Chinese Wu et al. (2025); Zhang et al. (2025); Xie et al. (2024); Nie et al. (2025). This asymmetry likely reflects the dominant role of English and other high-resource languages during pretraining and fine-tuning of multilingual language models.
@@ 5.3 Impact of Model Size and Capability
Figure 3: Generality scores for BLOOMZ and LLaMA models across increasing model sizes. ReMaKE Wang et al. (2024d) (BLOOMZ) results are extracted for mzsRE generality, averaged across 10 languages. IKE Nie et al. (2025) (LLaMA) results are extracted for mzsRE generality under zero-shot settings, averaged across 53 languages. Results show that generality consistently improves with model scale for both models.
@@ 6 Discussion
Multilingual knowledge editing has made progress, but major limitations remain. In this section, we identify persistent challenges and outline opportunities for future work. We organize these into thematic groups, reflecting both empirical trends and conceptual gaps in the literature.
@@ 6.1 Open Challenges
Multilingual LLMs encode languages into overlapping yet partially distinct internal subspaces Mousi et al. (2024), creating structural anisotropy that can hinder cross-lingual transfer. A key open problem for MKE is to bridge these spaces by using shared regions and handling language-specific pockets.
@@
Most MKE studies evaluate only high-resource languages (e.g., English, Chinese). Systematic evaluation on low-resource, morphologically complex, and under-represented languages (e.g., Swahili, Urdu) remains largely absent. In addition, current MKE benchmarks largely assume that factual knowledge is language-invariant. This overlooks culturally specific knowledge that may not have direct equivalents across languages, which is especially important for low-resource languages where such references are more prominent. Developing evaluation resources that account for cultural grounding remains an open challenge.
@@
Stability and Model Collapse
Another critical challenge, underexplored in multilingual settings, is model collapse, where even a few edits can substantially degrade an LLM’s general performance Green et al. (2025). Monolingual studies, Yang et al. (2024); Gupta et al. (2024) show that edits may trigger instability, and multilingual scenarios, with misaligned representational spaces, exacerbate this risk. Evaluating MKE methods should therefore consider both the success of edits and the preservation of model capabilities, particularly under sequential or batched updates where collapse risks accumulate.
Method Family Sequential Batch Notes
Parameter Editing
✗ (ROME)
✓ (LU-LAFNs)
$\sim$ (MEMLA)
✗ (ROME)
✓ (MEMIT, MEMLA, LU-LAFNs)
MEMIT supports batch;
LU-LAFNs allow partial
sequential edits.
Memory-Based
Edits stored externally;
supports dynamic updates
and reversibility.
Fine-Tuning
✗ (FT, LoRA-FT)
✓ (X-KDE, LTE)
✗ (FT, LoRA-FT)
$\sim$ (X-KDE, LTE)
High cost per update;
X-KDE enables reusable
batch edits.
Hypernetwork-Based
Meta-learned edits;
LiME supports low-latency,
language-conditioned updates.
Table 3: Compatibility of MKE method families with sequential and batch editing. ✓= supported; $\sim$ = partial; ✗= not supported.
Sequential and Batch Editing
Realistic deployment requires models to support not only isolated edits but also sequential updates and batched revisions. Most current MKE methods are not evaluated in these conditions. Memory-based and hypernetwork approaches (e.g., ReMaKE, LiME) naturally support sequential updates via externalized edits, while parameter editors such as MEMIT suit batch updates. Instruction-tuned methods (e.g., X-KDE) enable reusable edit behavior in batches but incur higher training costs. Table 3 summarizes relative compatibility across method families.
@@
The field still lacks unified evaluation protocols and benchmarks. While MLaKE and MZSRE are important steps forward, inconsistencies in construction, metrics, and coverage persist. Many rely on machine translation outputs, introducing noise. Low-resource and morphologically rich languages remain underrepresented.
Domain-specific Editing
Current MKE research is almost entirely evaluated on general-purpose factual knowledge. The impact of editing in specialized domains such as biomedical, legal, or culturally grounded knowledge remains unexplored. Developing benchmarks and methods for domain-specific MKE would broaden applicability and expose new challenges in precision and transfer.
@@ 6.2 Opportunities
Multi-source Editing
Most current MKE methods assume monolingual edits, with cross-lingual propagation only evaluated after the fact. A few approaches (e.g., LU-LAFNs, LiME) explicitly support multi-source editing by learning language-agnostic representations, allowing the same fact to be updated jointly across languages. This capability remains underexplored but represents an important opportunity for ensuring consistency and minimizing interference in multilingual models.
@@
6.3 Connections to Adjacent Research Areas
Hallucination mitigation
Knowledge editing has been explored as a way to correct hallucinations by directly updating model representations, with benchmarks such as HalluBench Huang et al. (2025) and TruthX Zhang et al. (2024b). However, evaluation remains English-only, and studies on multilingual hallucinations focus on measuring or reducing rates without using editing. Thus, the intersection of hallucination mitigation and MKE represents an exciting frontier for future research.
Mechanistic interpretability and cross-lingual transfer
Many editing methods, particularly parameter-based ones, build on mechanistic interpretability, assuming factual knowledge can be localized in neural circuits. Yet little is known about whether such circuits align across languages or how misalignment affects propagation. Exploring this link could ground MKE theoretically and explain why some edits generalize cross-lingually while others remain language-specific. Together, these connections show that MKE both informs and benefits from adjacent areas, though most current work remains monolingual.
Acknowledgments
We thank the anonymous reviewers and the meta-reviewer for their thoughtful and constructive feedback. Their comments helped us refine the scope of this survey, improve the clarity of our presentation, and integrate additional discussions (e.g., model collapse, sequential and batch editing, and connections to adjacent areas). We are grateful for their input, which has significantly strengthened the final version of this paper.
@@ Acknowledgments
Ethical Consideration
This work is a survey of existing research on multilingual knowledge editing and does not involve any new data collection, user studies, or model deployment. We do not foresee any direct ethical concerns. All referenced works have been appropriately cited, and any datasets or models discussed were originally released under appropriate terms by their respective authors. AI tools were used to rephrase and improve exposition of sections of the paper.
@@ Appendix A Details on MKE Benchmarks
MzSRE Wang et al. (2024d) is a machine-translated version of the ZsRE dataset in 12 languages: English, Czech, German, Dutch, Spanish, French, Portuguese, Russian, Thai, Turkish, Vietnamese, and Chinese. Translations were generated using Google Translate without human revision. Each sample is used to evaluate reliability, generality, locality, and portability.
@@
Eva-KELLM Wu et al. (2023) extends knowledge editing evaluation to full documents. Starting from the Counterfact dataset Meng et al. (2023a), ChatGPT was used to generate documents corresponding to each edited fact. The dataset consists of 8,882 English documents and 6,930 Chinese documents. [^4] Referred to as DocTer as of July 24, 2025
@@
BabelEdits  Green et al. (2025) is a knowledge editing benchmarking comprising 60 languages. It combines multlingual synsets from babel net with marker based translation to ensure entity (subject) translation quality. A comparison between the dataset is shown in Table 4